A circuit consists of a resistance R connected to n similar cells.If the current in the circuit is the same whether the cells are connected in series or in parallel,then the internal resistance of each cell is given by


a.r=R/n


b.r=nR


c.r=R


d.r=1/R

Grade:12th Pass

1 Answers

Sher Mohammad IIT Delhi
9 years ago
R=r+.........+r=nr
1/R=1/r+........+1/r=n/r
R=r/n=nr
imply n=1, r=R.

sher mohammad
b.tech, iit delhi